# Leveraging Social Determinants of Health in Alzheimer’s Research Using LLM-Augmented Literature Mining and Knowledge Graphs

### v1

@@ Introduction
As the most common type of dementia among the aging population globally, Alzheimer’s disease (AD) is a complex neurodegenerative disorder characterized by progressive cognitive decline and memory loss 1. The prevalence of AD and related dementias is projected to impact more than 100 million individuals worldwide by 2050 2. It presents critical challenges for patients and significant burdens for the public health system 3. To date, the underlying etiology of AD is still largely unknown, primarily due to the heterogeneity of the disease 4; 5; 6. It is widely believed that there is not a single cause of AD but rather a number of different factors that affect individuals differently 7. These factors include both medical and nonmedical ones, which can have a profound effect on a person’s risk. Since currently limited pharmacological treatments exist for AD 3, many previous studies have been attempting to identify and address risk factors that are modifiable to prevent or delay the disease progression.
@@
In this work, we present a novel framework of LLM-augmented literature mining to extract SDoH knowledge in the context of AD and integrate it with existing biomedical knowledge through knowledge graph construction and integration. We first collect relevant articles from PubMed and perform named entity recognition (NER) on the articles to identify biomedical entities and SDoH entities, utilizing OpenAI’s latest pretrained LLM GPT-4o 39 as well as classic NLP methods. For SDoH and biomedical entities, we then employ GPT’s language abilities to extract their potential relations and filter for co-occurrences to ensure accuracy, resulting in triplets of (biomedical entity, relation, SDoH entity) which form an AD-related SDoH knowledge graph. Furthermore, we integrate our SDoH knowledge graph with a biomedical subgraph derived from the general-purpose PrimeKG to tailor for AD-SDoH study. Finally, following previous work 40; 41; 42, we assess our new knowledge graph regarding the impact of SDoH on knowledge discovery in AD, by evaluating the performance of graph embedding-based link prediction tasks with graph neural networks 43. In summary, our main contributions can be listed as follows:
@@ Methods
Figure 1 presents the overview of our framework for generating an AD-related SDoH knowledge graph from PubMed literature. This process begins with the selected abstracts as described in the above section, and each of the abstract is subsequently processed through the following multi-step procedures. The detailed GPT-4o prompts with API calling and all other codes are available on our Github repo, due to the space limit in the paper.
@@ Results
As shown in Table 3, by merging SDoH information into subPrimeKG, SDoHenPKG shows a significant improvement in the link prediction task across all selected relations compared to subPrimeKG. This enhancement is evident from the consistently higher MRR values observed in SDoHenPKG, along with the p-values that indicate these improvements are statistically significant for most relations. This result suggests that the integration of SDoH data appears to provide additional context that helps the model better infer missing links related to AD.
Table 3: Results of the general AD link predictions task across seven relations. The relation names are abbreviated as follows: A for Anatomy to Gene/Protein, B for Disease to Drug, C for Disease to Effect, D for Disease to Gene/Protein, E for Drug to Effect, F for Biological Process to Gene/Protein, and G for Drug to Gene/Protein. The MRR values are the averages of five repeated experiments with standard deviations. * indicates significance levels from a two-tailed paired t-test (p-value $<$ 0.05).
Edges A B C D E F G
@@
Gene APOE TREM2 CD33 CLU BIN1 CR1 SORL1 CD2AP
@@
As we can see in Table 4, SDoHenPKG, which incorporates SDoH, generally outperformed subPrimeKG in predicting gene relationships, as demonstrated by higher MRR values across all the selected genes. The p-values indicate statistically significant improvements in predictive performance for several genes with SDoHenPKG. Specifically, genes like TREM2, BIN1, CR1, SPI1, INPP5D, ABI3, and C7 showed significant improvements, validating the integration of SDoH into the KG. This result strongly supports the benefit of incorporating SDoH data into etiology research, particularly for complex diseases like Alzheimer’s where biological and social factors are closely interlinked.
@@
For the top 1,000 (0.0001‰) predicted gene-gene edges with the highest probabilities, we conducted a PubMed search to check each gene pair’s co-occurrence and recorded the number of relevant articles. Notably, only 41 out of the 1,000 gene pairs had zero co-occurrence in the PubMed database, indicating that our predictions are both accurate and reasonable. Most gene pairs exhibited relatively low frequency, suggesting that they have not been extensively studied, thereby highlighting opportunities for future research. Interestingly, one gene pair, IL1B and TNF, had the highest co-occurrence frequency of 215,719 but was not included in PrimeKG. This finding illustrates that our prediction task can also be used to identify omissions or gaps in existing knowledge graphs. The complete list of all the predicted edges is available at our Github repo.
@@ Acknowledgments
This work was supported in part by the NIH grants U01 AG066833, U01 AG068057, and R01 AG071470.

### v2

@@ Introduction
As the most common type of dementia among the aging population globally, Alzheimer’s disease (AD) is a complex neurodegenerative disorder characterized by progressive cognitive decline and memory loss 1. The prevalence of AD and related dementias is projected to impact more than 100 million individuals worldwide by 2050 2. It presents critical challenges for patients and significant burdens for the public health system 3. To date, the underlying etiology of AD is still largely unknown, primarily due to the heterogeneity of the disease 4; 5; 6. It is widely believed that there is not a single cause of AD but rather a number of different factors that affect individuals differently 7. These factors include both medical and nonmedical ones, which can have a profound effect on a person’s risk for AD. Since currently limited pharmacological treatments exist for AD 3, many previous studies have been attempting to identify and address risk factors that are modifiable to prevent or delay the disease progression.
@@
In this work, we present a novel framework of LLM-augmented literature mining to extract SDoH knowledge in the context of AD and integrate it with existing biomedical knowledge through knowledge graph construction and integration. We first collect relevant articles from PubMed and perform named entity recognition (NER) on the articles to identify biomedical entities and SDoH entities, utilizing OpenAI’s latest pretrained LLM GPT-4o 39 as well as classic NLP methods. For SDoH and biomedical entities, we then employ GPT’s language abilities to extract their potential relations and filter for co-occurrences to ensure accuracy, resulting in triplets of (biomedical entity, relation, SDoH entity) which form an AD-related SDoH knowledge graph. Furthermore, we integrate our SDoH knowledge graph with a biomedical subgraph derived from the general-purpose PrimeKG to tailor for AD-SDoH study. Finally, following previous work 40; 41; 42, we assess our new knowledge graph regarding the impact of SDoH on knowledge discovery in AD, by evaluating the performance of graph embedding-based link prediction tasks with graph neural networks 43.
In summary, our main contributions can be listed as follows:
@@ Methods
Figure 1 presents the overview of our framework for generating an AD-related SDoH knowledge graph from PubMed literature. This process begins with the selected abstracts as described in the above section, and each of the abstract is subsequently processed through the following multi-step procedures. The detailed GPT-4o prompts with API calling, knowledge graph visualization, and all other codes are available on our Github repo, due to the space limit in the paper.
@@ Results
As shown in Table 3, by merging SDoH information into subPrimeKG, SDoHenPKG shows a significant improvement in the link prediction task across all selected relations compared to subPrimeKG. This enhancement is evident from the consistently higher MRR values observed in SDoHenPKG, along with the p-values that indicate these improvements are statistically significant for most relations. Significant improvements are observed in the Disease-Effect and Disease-Gene/Protein relations, indicating that SDoH factors may provide additional context for predicting disease progression and interactions. Similarly, improvements in Drug-Effect and Drug-Gene/Protein relations suggest that SDoH factors may impact drug efficacy, aligning with existing observations that SDoH can affect adherence, access, tolerance, and the overall clinical impact of dementia medications. While other relations show minor or insignificant improvements, the consistent pattern across key biomedical interactions highlights the relevance of SDoH in enhancing knowledge graph predictions. This result suggests that the integration of SDoH data appears to provide additional context that helps the model better infer missing links related to AD.
Table 3: Results of the general AD link predictions task across seven relations. The MRR values are the averages of five repeated experiments with standard deviations. * indicates significance levels from a two-tailed paired t-test (p-value $<$ 0.05).
Anatomy-Gene/Protein Disease-Drug Disease-Effect Disease-Gene/Protein Drug-Effect Biological Process-Gene/Protein Drug-Gene/Protein
@@
APOE TREM2 CD33 CLU BIN1 CR1 SORL1 CD2AP
@@
As we can see in Table 4, SDoHenPKG, which incorporates SDoH, generally outperformed subPrimeKG in predicting gene relationships, as demonstrated by higher MRR values across all the selected genes. The p-values indicate statistically significant improvements in predictive performance for several genes with SDoHenPKG. Specifically, genes like TREM2, BIN1, CR1, SPI1, INPP5D, ABI3, and C7 showed significant improvements, validating the integration of SDoH into the KG. This result shows that SDoH knowledge augments gene-gene link predictions, suggesting that social factors influence gene interactions relevant to AD pathology, which strongly supports the benefit of incorporating SDoH data into etiology research, particularly for complex diseases like Alzheimer’s where biological and social factors are closely interlinked.
@@
For the top 1,000 (0.0001‰) predicted gene-gene edges with the highest probabilities, we conducted a PubMed search to check each gene pair’s co-occurrence and recorded the number of relevant articles. Notably, only 41 out of the 1,000 gene pairs had zero co-occurrence in the PubMed database, indicating that our predictions are both accurate and reasonable. The overlap of predicted gene-gene edges with existing literature validates the reliability of our SDoH-augmented knowledge graph, demonstrating that the inclusion of SDoH information contributes to predicting biologically meaningful relationships rather than introducing arbitrary or random edges. Most gene pairs exhibited relatively low frequency, suggesting that they have not been extensively studied, thereby highlighting opportunities for future research. Interestingly, one gene pair, IL1B and TNF, had the highest co-occurrence frequency of 215,719 but was not included in PrimeKG. This finding illustrates that our prediction task can also be used to identify omissions or gaps in existing knowledge graphs. The complete list of all the predicted edges is available at our Github repo.
@@ Acknowledgments
This work was supported in part by the NIH grants P30 AG073105, U01 AG066833, U01 AG068057, and R01 AG071470.